H-atom is exposed to electromagnetic radiation of 1028 A and gives out induced radiations. Calculate the wavelenth of induced radiations.

Grade:11

1 Answers

Ramesh V
14 years ago

H-atom is exposed to UV radiation having wavelength 102.8nm

the energy of photon corresponding to 102.8 nm waveegth is

E = hc/λ = 1242 eV-nm /102.8 nm

   = 12.42 eV

The energy needed to take athe H atom frm ground state to first excited state is

                    E2-E1 - 13.6 eV - 3.4 eV = 10.2 eV

and to  the second excited state is

                    E3-E1 - 13.6 eV - 1.5 eV = 12.1 eV

and to  the third excited state is

                    E4-E1 - 13.6 eV - 0.85 eV = 12.75 eV

thus 12.42 eV photon has large probability of being absorbed in 12.1 eV

The corresponding  waveegth is: λ = 1242/12.1

                                                               =103 nm

---

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and
we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch